OpenStudy (anonymous):

5/12 divided 1/2 X 7/6= 35/144 is it right or where did i go wrong

hero (hero):

Are you asking if \[\frac{5}{12} \div \frac{1}{2} \times \frac{7}{6} = \frac{35}{144}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well in the problem i have their arent parentheses, just reads to simplify

hero (hero):

Okay

hero (hero):

So that means, all you have to do is flip the 1/2: \[\frac{5}{12} \times \frac{2}{1} \times \frac{7}{6}\]

hero (hero):

Now you can just multiply the numerators then the denominators.

OpenStudy (anonymous):

ok il do that

OpenStudy (anonymous):

35/36

hero (hero):

Actually, If you write it like this: \[\frac{(5)(2)(7)}{(12)(6)}\] You'll realize that you can cancel 2 and 6

hero (hero):

Good job

hero (hero):

\[\frac{(5)(7)}{(12)(3)} = \frac{35}{36}\]

hero (hero):

Remember that when you're dividing two fractions you have to flip the second fraction according to the reciprocal rule.

OpenStudy (anonymous):

ok thanks for that i was confused as to why you flipped it/... thank u

OpenStudy (anonymous):

when it is a for example 36/84 if you cant divide after that you find the lcm to get an answer ? this question also refers to when dividing

hero (hero):

You probably forgot to reduce factors of one before multiplying

OpenStudy (anonymous):

look 4/3 divided by 9/28 ( i already flipped the 2nd part) = 36/84 lcm is 12 my final answer would be 3/7

hero (hero):

I would do it this way: \[\frac{4}{3} \div\frac{9}{28} = \frac{4}{3} \times \frac{28}{9} = \frac{(4)(4)(7)}{(3)(3)(3)} = \] Doing it this way we see that nothing cancels. So we have to multiply it all out.

OpenStudy (anonymous):

ok got it

hero (hero):

However, after multiplying, you'll have an improper fraction.

hero (hero):

You'll have to re-write it as a whole number plus a fraction